# High-cadence observations of galactic nuclei by the future two-band UV-photometry mission QUVIK

### v1

@@ Abstract
The Quick Ultra-Violet Kilonova surveyor (QUVIK), a two-band UV space telescope approved for funding as a Czech national science and technology mission, will focus on detecting early UV light of kilonovae (Werner et al., 2024). In addition, it will study the UV emission of stars and stellar systems (Krtička et al., 2024) as well as the intense and variable emission of active galactic nuclei (AGN) or galactic nuclei activated by tidal disruption events (Zajaček et al., 2024). In this contribution, we describe the role of this small ($\sim 30$-cm diameter) UV telescope for studying bright, nearby AGN. With its NUV and FUV bands, the telescope will perform high-cadence ($\sim 0.1$-1 day) two-band photometric monitoring of nearby AGN ($z<1$), which will allow us to probe accretion disk sizes/temperature profiles via photometric reverberation mapping. Thanks to its versatility, QUVIK will be able to perform a moderately fast repointing ($<20$ min) to target candidates for tidal disruption events (TDEs). Early detection of the UV emission following a TDE optical flare, in combination with the subsequent two-band UV monitoring performed simultaneously with other observatories, will enable us to infer the time delay (or its lack of) between the optical, UV, and X-ray emission. In combination with theoretical models, it will be possible to shed more light on the origin of the UV/optical emission of TDEs. Furthermore, the two-band monitoring of nuclear transients will be beneficial in distinguishing between TDEs (nearly constant blue colour) and supernovae (progressive reddening).
@@ 1 Introduction
The UV light consists of both continuum and line emission contributions. The continuum can be of thermal or non-thermal origin, which eventually determines the slope of the continuum spectral energy distribution (SED) in the UV domain or, in other words, the UV colour between far-UV (FUV) and near-UV (NUV) bands. Assuming optically thick, blackbody emission, in the FUV domain ($\sim 150$ nm) objects with the effective temperature of $T_{\rm NUV}\simeq b/\lambda=19320\,(b/0.29\,{\rm cm\,K})(\lambda/150\,{\rm nm})^{-1}\,{\rm K}$ emit the most (Wien displacement’s law), while in the NUV domain ($\sim 300$ nm) the effective temperature drops by a factor of two to $T_{\rm FUV}\sim 9660\,{\rm K}$. In terms of stellar spectral types, these temperatures correspond to the spectral types B (10 00030 000similar-toabsent1000030000$\sim$ 10\,000-30\,000∼ 10 000 - 30 000 K) and A (7 30010 000similar-toabsent730010000$\sim$ 7\,300-10\,000∼ 7 300 - 10 000 K). In general, different types of celestial objects and associated processes can be studied in both NUV and FUV bands:
  • rapid stellar transients, such as the UV emission of kilonovae (depending on the explosion scenario), $\gamma$-ray burst afterglows, and supernovae of type I and II[2],
@@
Many significant measurements in the UV domain, including those related to the variability and spectral properties of AGN, were performed by the International Ultraviolet Explorer (IUE), which was launched in 1978. The mission equipped with a mirror of 45 cm lasted for 18 years until 1996, and it was capable of short-wavelength (115-200 nm) and long-wavelength (185-330 nm) UV spectroscopy [6]. It provided several key measurements, including the studies of stellar winds and interstellar medium extinction, as well as detailed observations of the supernova SN 1987A, which deviated from standard stellar evolution models at the time. Concerning AGN, IUE performed detailed monitoring of the nearby bright Seyfert galaxy NGC 4151 ($z=0.0033)$, which exhibited variability on timescales of only several days [7]. More importantly, with respect to the continuum variability of NGC 4151, IUE detected time delays of several days for the UV broad lines (CIV, HeII, CIII], and MgII)[8], thus confirming the reprocessing scenario of the broad-line region (BLR) clouds that surround the supermassive black hole (SMBH) at a specific distance range that is traced by the broad lines of a certain ionization potential[9]. In combination with the detected line dispersion, it was possible to infer the virial SMBH mass of $M_{\bullet}=(4.14\pm 0.74)\times 10^{7}\,M_{\odot}$ using $M_{\bullet}=f_{\sigma}c\tau_{\rm BLR}\sigma^{2}/G$ where $f_{\sigma}$ is the virial factor, $\tau_{\rm BLR}$ is the time delay of the line variability, and $\sigma$ is the line dispersion [8], with $c$ and $G$ as the speed of light and the gravitational constant. The monitoring of NGC 5548 ($z=0.01651$) by IUE led to the determination of the spatial offset of the UV FeII/Balmer continuum complex (“small blue bump”) between 220 and 420 nm from the SMBH ($\sim 10$ days, which corresponds to 1 700similar-toabsent1700$\sim$ 1\,700∼ 1 700 AU)[10]. The UV FeII pseudocontinuum/Balmer continuum was found to be variable with an amplitude of $\sim 20\%$. The UV FeII emission, consisting of many FeII line transitions is an important contributor to the line cooling of the circumnuclear gas.
@@
Furthermore, important UV observations are currently provided by the Hubble Space Telescope (HST) launched in 1990 and equipped with a 2.4-meter mirror, which directs the incoming photons into several near-UV and far-UV imagers and spectrographs, in particular the Advanced Camera for Surveys (ACS; near-UV imaging), Cosmic Origins Spectrograph (COS; UV spectra in the range 115-320 nm), Space Telescope Imaging Spectrograph (spatially resolved spectroscopy from 115 nm to 1030 nm, with the high spatial resolution echelle spectroscopy in the UV domain and photon-tagging to achieve a high temporal resolution), and Wide Field Camera 3 equipped with a UVIS channel for imaging between 200 and 1000 nm. Another relevant instrument is the Neil Gehrels Swift Observatory launched in 2004, which is equipped with an Ultra-Violet/Optical Telescope (UVOT) instrument onboard, in addition to $\gamma$-ray and X-ray sensitive detectors. Hence, Swift is capable of intense multiwavelength monitoring of all kinds of transients, from $\gamma$-ray bursts to galactic nuclei transients. The UVOT telescope is a modified Ritchey-Chrétien with a 30-cm mirror, $2.5^{\prime\prime}$ point spread function (PSF; at 350 nm), and a sensitivity of 22.3 AB magnitude (in 1000 seconds in the B band). UVOT can perform UV imaging and photometry in four UV channels (B, UVW1, UVM2, and UVW2) from $\sim 150$ nm to $\sim 390$ nm[15]. Concerning AGN research, HST and Swift provided high-cadence ($\sim 1$ day) UV light curves for accretion-disk reverberation-mapping (RM) of nearby type 1 AGN (for instance, NGC 5548 at $z=0.01651$[16]). The interband time delays as a function of the band central wavelength are generally consistent with the standard accretion disk predictions for the reprocessing of the continuum emission, i.e. the relation $\tau(\lambda)=\alpha[(\lambda/\lambda_{0})^{\beta}-1]$ with $\beta\simeq 4/3$ within the uncertainties, where $\tau$ is the time lag of the continuum emission at the wavelength $\lambda$ expressed with respect to the wavelength $\lambda_{0}$ and $\beta$ is the slope related to the temperature profile of the accretion disk, $T_{\rm disk}\propto r^{-1/\beta}$, and $\alpha$ is the normalization coefficient. Two basic discrepancies were found with the predictions of the standard accretion disk theory: (i) the time delays/accretion-disk sizes tend to be larger by a factor of $\sim 3$, (ii) the $U$-band time delays are significantly larger than the best-fit time-delay/wavelength relation due to the contamination by the broad-line emission contribution (contamination by the Balmer continuum in the $U$ band reaches $\sim 20\%$)[17]. Among other galactic nucleus discoveries, the Swift Observatory target-of-opportunity monitoring recently led to the discovery of repeating nuclear transients with periods of about one month, which is intermediate between the X-ray quasi-periodic erupters (a few days) and the repeating partial TDEs (hundreds of days), see in particular the study of Swift J0230-28 with the periodicity of $\sim 22$ days[18, 19, 20].
@@
AstroSat was launched in 2015 with the aim to carry out multiwavelength observations and imaging of variable sources on different scales, from stars and stellar binaries, to galaxies, galactic nuclei, and galaxy groups and clusters. It performs observations in the optical, UV, soft, and hard X-ray domains. Its Ultra-Violet Imaging Telescope (UVIT) with a mirror size of $\sim 40$ cm can perform imaging with a field of view of $\sim 28^{\prime}$ in diameter in three spectral channels, 130-180 nm, 180-300 nm, and 320-530 nm, with the possibility of performing a slitless grating low-resolution spectroscopy ($R\sim 100$). A relatively high angular resolution in the UV domain $\sim 1.8^{\prime\prime}$ allows the decomposition of a galaxy surface brightness into the host and the point-source AGN. AstroSat observations in the X-ray/UV domain led to tighter constraints of broad-band AGN spectra. In a few cases, relatively large accretion-disk inner radii of $\sim 35$-150 gravitational radii ($r_{\rm{g}}$) were found[21, 22], which may correspond to the transition between the outer standard thin disk and the inner hot advection-dominated accretion flow[23] for lower-accreting AGN.
@@
The Quick Ultra-Violet Kilonova surveyor, QUVIK mission is designed to provide dedicated rapid follow-up observations of transients discovered by ULTRASAT and other wide-field sky surveys [2, 3, 4]. QUVIK has been selected as a Czech national science and technology mission to be launched towards the end of this decade (2029-2030). The mission will provide sensitive photometry in NUV and FUV bands complementary to the bandpass of ULTRASAT. The mission baseline includes a NUV band of 260–360 nm and a FUV band of 140–190 nm [2]. Importantly, in the two UV bands, the mission will provide simultaneous photometry, which will also be important for the study of AGN [4], in particular for the reverberation mapping of accretion flows. In addition, the simultaneous monitoring in the two UV bands will constrain the FUV-NUV colour evolution of transients, which will be crucial for their classification. In particular it will help distinguish a TDEs (constant blue colour) from a supernovae (progressive reddening). In the NUV band, the field of view is aimed at $\gtrsim 1^{\circ}\times 1^{\circ}$ with a point spread function FWHM of $\lesssim 2.5$ arcseconds, which will allow to isolate point sources (AGN) from the extended background emission (host galaxy).The primary objective of QUVIK is to perform a rapid follow-up of transient sources with the observation start latency of $\lesssim 20$ minutes after receiving a trigger from wide field-of-view surveys (e.g., ULTRASAT, Legacy Survey of Space and Time by the Vera C. Rubin observatory) and gravitational-wave observatories detecting potential neutron star – neutron star mergers (see Section 2 for more details). Furthermore, QUVIK will be versatile enough to increase the monitoring cadence of transients according to the need, going potentially down to $\sim 0.1$ days. In addition, complementary to the ULTRASAT UV survey mission, QUVIK will conduct high-cadence monitoring of selected targets in two UV bands, which will, in synergy with ULTRASAT, result in constraints on the temporal evolution of the UV spectral slopes determined by the two-band flux density measurement by QUVIK and the one-band measurement by ULTRASAT.
QUVIK will be launched to a sun-synchronous low-Earth orbit at an altitude of $\sim 550\,{\rm km}$ with an orbital period of $\sim 90$ minutes. In Fig. 1, we show an artistic image of a two-band UV mission analogous to the QUVIK featuring the main scientific aims of such a mission: stellar transients (kilonovae, supernovae, novae), stars and stellar populations, and AGN including nuclear transients.
@@
This contribution is structured as follows. In Section 2, we introduce the primary scientific objective of QUVIK – the early detection of the UV light of kilonovae. In Section 3, we feature one of the secondary scientific objectives – AGN and (repeating) nuclear transients. In Section 4, we briefly outline the synergy of QUVIK monitoring and observations with other missions available at the time of its launch and active phase (2029-2032). We summarize the main characteristics of the QUVIK mission in Section 5.
@@ 3 Exemplary secondary objective: Active galactic nuclei and nuclear transients
The QUVIK mission will have the ability to quickly observe targets of opportunity and provide rapid observations in NUV and FUV bands. This is ideal to study many types of transient events beyond the primary scientific objective, such as TDEs, AGN flaring events, changing look AGN (CL-AGN), and microlensing high magnification events (HMEs). Furthermore, it offers the possibility to perform reverberation mapping of the AGN’s inner accretion disk due to the NUV and FUV bands.
@@
where $M_{\bullet}$ is the SMBH mass, $G$ is the gravitational constant, $m_{\rm p}$ is the proton mass, $\sigma_{\rm T}$ is the Thomson scattering cross-section, and $c$ is the speed of light. Thus quasars typically accrete of the order of one Solar mass per year. It is convenient to express the accretion rate of AGN relative to the Eddington rate. Assuming the same radiative efficiency, this is referred to as the relative accretion rate, $\dot{m}\equiv\dot{M}/\dot{M}_{\rm Edd}$, or the Eddington ratio (often denoted as $\lambda_{\rm Edd}=\dot{m}$). The standard thin-disk spectrum is given as the sum of black-body spectra of thin rings with temperatures following the power-law profile $T_{\rm disk}\propto r^{-3/4}$ as a function of radius. In Fig. 2, we show exemplary spectral energy distributions (SEDs) of standard accretion disks surrounding the SMBHs of $M_{\bullet}=10^{8}\,M_{\odot}$ and $10^{9}\,M_{\odot}$, whose relative accretion rates are $\dot{m}=0.1$ and they are viewed at $\iota=20^{\circ}$. We see the basic trend that SMBHs with higher masses have larger UV luminosities and their SEDs are softer, i.e. the SED peak is shifted towards the NUV waveband. When unobscured by intervening circumnuclear and host dust, AGN of type I can be detected by the QUVIK UV mission across a wide redshift range from the local universe to the intermediate redshifts around $z\sim 1$, see Fig. 3. This is especially the case of AGN with SMBHs of a higher mass of $M_{\bullet}=10^{8}-10^{9}\,M_{\odot}$. For a relative accretion rate of $\dot{m}=0.1$, the inclination of $\iota=20^{\circ}$, and the SMBH mass of $M_{\bullet}=10^{8}\,M_{\odot}$, the estimated UV AB magnitude is $m_{\rm AB}=10.7$ mag at $\lambda=310$ nm and $z=0.01$. At $z=1$, it is $m_{\rm AB}=21.6$ mag, which is at the sensitivity limit of the QUVIK in the NUV band. However, for larger SMBH masses and higher relative accretion rates, the AGN brightness in the UV domain increases, hence the prospects for the detection of type I AGN and quasars at low- to intermediate-redshift are good (see Fig. 3 for the comparison between $M_{\bullet}=10^{8}\,M_{\odot}$ and $M_{\bullet}=10^{9}\,M_{\odot}$ at different redshifts).

### v2

@@ Abstract
The Quick Ultra-VIolet Kilonova surveyor (QUVIK), a two-band UV space telescope approved for funding as a Czech national science and technology mission, will focus on detecting early UV light of kilonovae (Werner et al., 2024). In addition, it will study the UV emission of stars and stellar systems (Krtička et al., 2024) as well as the intense and variable emission of active galactic nuclei (AGN) or galactic nuclei activated by tidal disruption events (Zajaček et al., 2024). In this contribution, we describe the role of this small ($\sim 30$-cm diameter) UV telescope for studying bright, nearby AGN. With its NUV and FUV bands, the telescope will perform high-cadence ($\sim 0.1$-1 day) two-band photometric monitoring of nearby AGN ($z<1$), which will allow us to probe accretion disk sizes/temperature profiles via photometric reverberation mapping. Thanks to its versatility, QUVIK will be able to perform a moderately fast repointing (<20<20 min) to target candidates for tidal disruption events (TDEs). Early detection of the UV emission following a TDE optical flare, in combination with the subsequent two-band UV monitoring performed simultaneously with other observatories, will enable us to infer the time delay (or its lack of) between the optical, UV, and X-ray emission. In combination with theoretical models, it will be possible to shed more light on the origin of the UV/optical emission of TDEs. Furthermore, the two-band monitoring of nuclear transients will be beneficial in distinguishing between TDEs (nearly constant blue colour) and supernovae (progressive reddening) in the era of intensive wide-field surveys.
@@ 1 Introduction
The UV light consists of both continuum and line emission contributions. The continuum can be of thermal or non-thermal origin, which eventually determines the slope of the continuum spectral energy distribution (SED) in the UV domain or, in other words, the UV colour between far-UV (FUV) and near-UV (NUV) bands. Assuming optically thick, blackbody emission, in the FUV domain ($\sim 150$ nm) objects with the effective temperature of $T_{\rm NUV}\simeq b/\lambda=19320\,(b/0.29\,{\rm cm\,K})(\lambda/150\,{\rm nm})^{-1}\,{\rm K}$ emit the most (Wien displacement’s law), while in the NUV domain ($\sim 300$ nm) the effective temperature drops by a factor of two to $T_{\rm FUV}\sim 9660\,{\rm K}$. In terms of stellar spectral types, these temperatures correspond to the spectral types B ($\sim 10\,000-30\,000$ K) and A ($\sim 7\,300-10\,000$ K). In general, different types of celestial objects and associated processes can be studied in both NUV and FUV bands:
  • rapid stellar transients, such as the UV emission of kilonovae (depending on the explosion scenario), $\gamma$-ray burst afterglows, and supernovae of type I and II[2, 3],
@@
Many significant measurements in the UV domain, including those related to the variability and spectral properties of AGN, were performed by the International Ultraviolet Explorer (IUE), which was launched in 1978. The mission equipped with a mirror of 45 cm lasted for 18 years until 1996, and it was capable of short-wavelength (115-200 nm) and long-wavelength (185-330 nm) UV spectroscopy [7]. It provided several key measurements, including the studies of stellar winds and interstellar medium extinction, as well as detailed observations of the supernova SN 1987A, which deviated from standard stellar evolution models at the time. Concerning AGN, IUE performed detailed monitoring of the nearby bright Seyfert galaxy NGC 4151 ($z=0.0033)$, which exhibited variability on timescales of only several days [8]. More importantly, with respect to the continuum variability of NGC 4151, IUE detected time delays of several days for the UV broad lines (CIV, HeII, CIII], and MgII)[9], thus confirming the reprocessing scenario of the broad-line region (BLR) clouds that surround the supermassive black hole (SMBH) at a specific distance range that is traced by the broad lines of a certain ionization potential[10]. In combination with the detected line dispersion, it was possible to infer the virial SMBH mass of $M_{\bullet}=(4.14\pm 0.74)\times 10^{7}\,M_{\odot}$ using $M_{\bullet}=f_{\sigma}c\tau_{\rm BLR}\sigma^{2}/G$ where $f_{\sigma}$ is the virial factor, $\tau_{\rm BLR}$ is the time delay of the line variability, and $\sigma$ is the line dispersion [9], with $c$ and $G$ as the speed of light and the gravitational constant. The monitoring of NGC 5548 ($z=0.01651$) by IUE led to the determination of the spatial offset of the UV FeII/Balmer continuum complex (“small blue bump”) between 220 and 420 nm from the SMBH ($\sim 10$ days, which corresponds to $\sim 1\,700$ AU)[11]. The UV FeII pseudocontinuum/Balmer continuum was found to be variable with an amplitude of $\sim 20\%$. The UV FeII emission, consisting of many FeII line transitions is an important contributor to the line cooling of the circumnuclear gas.
@@
Furthermore, important UV observations are currently provided by the Hubble Space Telescope (HST) launched in 1990 and equipped with a 2.4-meter mirror, which directs the incoming photons into several near-UV and far-UV imagers and spectrographs, in particular the Advanced Camera for Surveys (ACS; near-UV imaging), Cosmic Origins Spectrograph (COS; UV spectra in the range 115-320 nm), Space Telescope Imaging Spectrograph (spatially resolved spectroscopy from 115 nm to 1030 nm, with the high spatial resolution echelle spectroscopy in the UV domain and photon-tagging to achieve a high temporal resolution), and Wide Field Camera 3 equipped with a UVIS channel for imaging between 200 and 1000 nm.
@@
Another relevant instrument is the Neil Gehrels Swift Observatory launched in 2004, which is equipped with an Ultra-Violet/Optical Telescope (UVOT) instrument onboard, in addition to $\gamma$-ray and X-ray sensitive detectors. Hence, Swift is capable of intense multiwavelength monitoring of all kinds of transients, from $\gamma$-ray bursts to galactic nuclei transients. The UVOT telescope is a modified Ritchey-Chrétien with a 30-cm mirror, $2.5^{\prime\prime}$ point spread function (PSF; at 350 nm), and a sensitivity of 22.3 AB magnitude (in 1000 seconds in the B band). UVOT can perform imaging and photometry in six UV/optical filters (UVW2, UVM2, UVW1, U, B, V) from $170$ to $650$ nm [16] within a $\times17$’ field. Concerning AGN research, both the HST and Swift provided high-cadence ($\sim 1$ day) UV light curves for accretion-disk reverberation-mapping (RM) of nearby type 1 AGN (for instance, NGC 5548 at z=0.01651z=0.01651[17]). The interband time delays as a function of the band central wavelength are generally consistent with the standard accretion disk predictions for the reprocessing of the continuum emission, i.e. the relation $\tau(\lambda)=\alpha[(\lambda/\lambda_{0})^{\beta}-1]$ with $\beta\simeq 4/3$ within the uncertainties, where $\tau$ is the time lag of the continuum emission at the wavelength $\lambda$ expressed with respect to the wavelength $\lambda_{0}$ and $\beta$ is the slope related to the temperature profile of the accretion disk, $T_{\rm disk}\propto r^{-1/\beta}$, and $\alpha$ is the normalization coefficient. Two basic discrepancies were found with the predictions of the standard accretion disk theory:
  • (i)
    the time delays/accretion-disk sizes tend to be larger by a factor of $\sim 3$,
  • (ii)
    the $U$-band time delays are significantly larger than the best-fit time-delay/wavelength relation due to the contamination by the broad-line emission contribution (contamination by the Balmer continuum in the $U$ band reaches $\sim 20\%$)[18]. In fact, the $U$-band excess is commonly detected. The HST spectroscopic monitoring of NGC 4593 showed that the feature is rather broad, extending from 300 to 400 nm, with a discontinuity at the Balmer jump (364.6 nm), which is consistent with the BLR contribution scenario [19].
Among other galactic nucleus discoveries, the Swift Observatory target-of-opportunity monitoring recently led to the discovery of repeating nuclear transients with periods of about one month, which is intermediate between the X-ray quasi-periodic erupters (QPEs; from a few hours to a day) and the repeating partial TDEs (hundreds of days), see in particular the study of Swift J0230-28 with the periodicity of $\sim 22$ days[20, 21, 22] and the “Ansky” source with the QPE period of $\sim 4.5$ days (ZTF19acnskyy[23, 24]).
@@
AstroSat was launched in 2015 with the aim to carry out multiwavelength observations and imaging of variable sources on different scales, from stars and stellar binaries, to galaxies, galactic nuclei, and galaxy groups and clusters. It performs observations in the optical, UV, soft, and hard X-ray domains. Its Ultra-Violet Imaging Telescope (UVIT) with a mirror size of $\sim 40$ cm can perform imaging with a field of view of $\sim 28^{\prime}$ in diameter in three spectral channels, 130-180 nm, 180-300 nm, and 320-530 nm, with the possibility of performing a slitless grating low-resolution spectroscopy ($R\sim 100$). A relatively high angular resolution in the UV domain $\sim 1.8^{\prime\prime}$ allows the decomposition of a galaxy surface brightness into the host and the point-source AGN. AstroSat observations in the X-ray/UV domain led to tighter constraints on broadband AGN spectra. In a few cases, relatively large accretion-disk inner radii of $\sim 35$-150 gravitational radii ($r_{\rm{g}}$) were found[25, 26], which may correspond to the transition between the outer standard thin disk and the inner hot advection-dominated accretion flow[27] for lower-accreting AGN.
@@
This contribution is structured as follows. We begin with Section 2 where we introduce the concept of the Czech ambitious space telescope QUVIK. In Section 3, we introduce the primary scientific objective of QUVIK – the early detection of the UV light of kilonovae. In Section 4, we feature one of the secondary scientific objectives – AGN and (repeating) nuclear transients. In Section 5, we briefly outline the synergy of QUVIK monitoring and observations with other missions available at the time of its launch and active phase (2029-2032). We summarize the main characteristics of the QUVIK mission in Section 6.
2 QUVIK Space Telescope – Czech Ambitious Space Mission
@@
The Quick Ultra-VIolet Kilonova surveyor, QUVIK mission [30] is designed to provide dedicated rapid follow-up observations of transients discovered by ULTRASAT and other wide-field sky surveys [3, 4, 5]. QUVIK has been selected as a Czech national science and technology mission to be launched towards the end of this decade (2029-2030). The satellite will operate on a Sun-synchronous low-Earth orbit with an orbital period of $\sim 90$ minutes. The primary mission shall last for three years. The mission will provide sensitive photometry in NUV and FUV bands complementary to the bandpass of ULTRASAT. The mission baseline includes a NUV band of 260–360 nm and a FUV band of 140–190 nm [3]. Importantly, in the two UV bands, the mission will provide simultaneous photometry, which will also be important for the study of AGN [5], in particular for the reverberation mapping of accretion flows. In addition, the simultaneous monitoring in the two UV bands will constrain the FUV-NUV colour evolution of transients, which will be crucial for their classification. In particular, it will help distinguish TDEs (constant blue colour) from supernovae (progressive reddening). In the NUV band, the field of view is aimed at $\gtrsim 1^{\circ}\times 1^{\circ}$ with a point spread function FWHM of $\lesssim 2.5$ arcseconds, which will allow isolating point sources (AGN) from the extended background emission (host galaxy).The primary objective of QUVIK is to perform a rapid follow-up of transient sources with the observation start latency of $\lesssim 20$ minutes after receiving a trigger from wide field-of-view surveys (e.g., ULTRASAT, Legacy Survey of Space and Time by the Vera C. Rubin observatory) and gravitational-wave observatories detecting potential neutron star – neutron star mergers (see Section 3 for more details). Furthermore, QUVIK will be versatile enough to increase the monitoring cadence of transients according to the need, going potentially down to $\sim 0.1$ days for bright enough sources ($m_{\rm AB}\lesssim 20$ mag), for which the total integration time will be $\lesssim 10$ minutes. This will enable to capture short-term UV variability on the timescale of a few hours. In addition, complementary to the ULTRASAT UV survey mission, QUVIK will conduct high-cadence monitoring of selected targets in two UV bands, which will, in synergy with ULTRASAT, result in constraints on the temporal evolution of the UV spectral slopes determined by the two-band flux density measurement by QUVIK and the one-band measurement by ULTRASAT. In Fig. 1, we show an artistic image of a two-band UV mission analogous to the QUVIK featuring the main scientific aims of such a mission: stellar transients (kilonovae, supernovae, novae), stars and stellar populations, and AGN, including nuclear transients. QUVIK will overall be constructed in a small-satellite design with the dimensions $70\times 70\times 110$ cm and the projected weight of $\sim 200$ kg. It will be equipped with the near-real-time communication for the transient alert reception and the moderately fast repointing capability with the observation start latency of $\lesssim 20$ minutes[3]. Its primary payload will be the Cassegrain telescope with a mirror diameter of $\sim 30$ cm. The NUV detector aims at a photometric sensitivity of 21.5 AB magnitude with a signal-to-noise ratio of 5 in one orbit. The overall simple and yet versatile concept of the QUVIK is in line with the 2020 US Decadal survey, which highlighted the role of small- and medium-sized missions in advancing the time-domain and multi-messenger astrophysics[31].
@@ 4 Exemplary secondary objective: Active galactic nuclei and nuclear transients
The QUVIK mission will have the ability to quickly observe targets of opportunity and provide rapid observations in NUV and FUV bands. This is ideal to study many types of transient events beyond the primary scientific objective, such as TDEs, AGN flaring events, changing look AGN (CL-AGN), and microlensing high-magnification events (HMEs). Furthermore, it offers the possibility to perform reverberation mapping of the AGN’s inner accretion disk due to the NUV and FUV bands.
@@
where $M_{\bullet}$ is the SMBH mass, $G$ is the gravitational constant, $m_{\rm p}$ is the proton mass, $\sigma_{\rm T}$ is the Thomson scattering cross-section, and $c$ is the speed of light. Thus quasars typically accrete of the order of one Solar mass per year. It is convenient to express the accretion rate of AGN relative to the Eddington rate. Assuming the same radiative efficiency, this is referred to as the relative accretion rate, $\dot{m}\equiv\dot{M}/\dot{M}_{\rm Edd}$, or the Eddington ratio (often denoted as $\lambda_{\rm Edd}=\dot{m}$). The standard thin-disk spectrum is given as the sum of black-body spectra of thin rings with temperatures following the power-law profile $T_{\rm disk}\propto r^{-3/4}$ as a function of radius. In Fig. 2, we show example spectral energy distributions (SEDs) of standard accretion disks surrounding the SMBHs of $M_{\bullet}=10^{8}\,M_{\odot}$ and $10^{9}\,M_{\odot}$, whose relative accretion rates are $\dot{m}=0.1$ and they are viewed at $\iota=20^{\circ}$. We see the basic trend that SMBHs with higher masses have larger UV luminosities and their SEDs are softer, i.e. the SED peak is shifted towards the NUV waveband. When unobscured by intervening circumnuclear gas and warm dust, AGN of type I can be detected by the QUVIK UV mission across a wide redshift range from the local universe to the intermediate redshifts around $z\sim 1$, see Fig. 3. This is especially the case of AGN with SMBHs of a higher mass of $M_{\bullet}=10^{8}-10^{9}\,M_{\odot}$. For a relative accretion rate of $\dot{m}=0.1$, the inclination of $\iota=20^{\circ}$, and the SMBH mass of $M_{\bullet}=10^{8}\,M_{\odot}$, the estimated UV AB magnitude is $m_{\rm AB}=10.7$ mag at $\lambda=310$ nm and $z=0.01$. At $z=1$, it is $m_{\rm AB}=21.6$ mag, which is at the sensitivity limit of the QUVIK in the NUV band. However, for larger SMBH masses and higher relative accretion rates, the AGN brightness in the UV domain increases, hence the prospects for the detection of type I AGN and quasars at low- to intermediate-redshift are favourable (see Fig. 3 for the comparison between $M_{\bullet}=10^{8}\,M_{\odot}$ and $M_{\bullet}=10^{9}\,M_{\odot}$ at different redshifts).